# Measurement-aligned Flow for Inverse Problem

### v1

@@ B.1 Addressing Gaussian noise
Consider noisy image restoration problems in the form of $y=Hx+\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\epsilon_{y}}$, where $\epsilon_{y}$ is the added noise. Then the measurement $y$ can be decomposed to the sum of clean measurement $y^{\text{clean}}\coloneqq Hx$ and measurement noise $\epsilon_{y}$. Calculating posterior mean $\tilde{x}_{0}^{\text{pm}}$ leads to:
$\tilde{x}_{0}^{\text{pm}}$ $=Y^{-1}[\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (43)
$=\hat{x}_{0}^{\theta}+(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+Y^{-1}H^{T}W^{-1}y$ (44)
where ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{T}W^{-1}\epsilon_{y}}$ is the extra noise introduced into $\tilde{x}_{0}^{\text{pm}}$ and will be further introduced into $x_{t-\Delta t}$.
To address Gaussian noise with known variance, we modify Eq. 12 and Eq. 3 as:
$\tilde{x}_{0}^{\text{pm}}=\hat{x}_{0}^{\theta}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}[(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (45)
@@
Then $\tilde{x}_{0}^{\text{pm}}$ is:
$\tilde{x}_{0}^{\text{pm}}$ $=\hat{x}_{0}^{\theta}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}[(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (47)
$=\underbrace{\hat{x}_{0}^{\theta}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+Y^{-1}H^{T}W^{-1}y^{\text{clean}}}_{\coloneqq\;\tilde{x}_{0}^{\text{clean}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ (49)
@@
$x_{t-\Delta t}$ $=a_{t}\tilde{x}_{0}^{\text{pm}}(t,x,y)+b_{t}x_{t}+\epsilon_{\text{new}},\quad\epsilon_{\text{new}}\sim\mathcal{N}(0,\Phi_{t})$ (50)
@@
Suppose $\Sigma_{t}=V\mathrm{diag}\{\lambda_{t1},\cdots,\lambda_{td}\}V^{T}$ $\Phi_{t}=Vdiag\{\gamma_{t1},\cdots\gamma_{td}\}V^{T}$. Then the introduced noise $\epsilon_{\text{intro}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{t}\sigma_{y}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ is still a Gaussian distribution: $\epsilon_{\text{intro}}\sim\mathcal{N}(0,VD_{t}V^{T})$, with $D_{t}=diag\{d_{t1},\cdots,d_{td}\}$:
@@
The choice of and $\Phi_{t}$ need to ensure the total noise injected to $x_{t-\Delta t}$ conforms the iteration in Eq. 3.
@@
Now we can yield $\epsilon_{\text{new}}$ by sampling from $\mathcal{N}(0,V\Gamma_{t}V^{T})$ to ensure that $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. We need to make sure $\lambda_{ti}$ guaratees the noise level of the introduced noise do not exceed the pre-defined noise level $c_{t}$, we also hope $\lambda_{t_{i}}$ as close as $1$ as possible. Therefore,
@@
B.2 Addressing unknown noise and non-differentiable measurements
As the measurement noise $\epsilon_{y}$ is non-Gaussian or unknown, it’s difficult to ensure the total noise introduced in $x_{t-\Delta t}$ is $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. In this case, we calculate $\tilde{x}_{0}^{\text{pm}}$ using Eq. 12 and update $x_{t-\Delta t}$ using Eq. 3. Then $\tilde{x}_{0}^{\text{pm}}$ is:
$\tilde{x}_{0}^{\text{pm}}$ $=\hat{x}_{0}^{\theta}+[(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (56)
$=\underbrace{\hat{x}_{0}^{\theta}+(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+Y^{-1}H^{T}W^{-1}y^{\text{clean}}}_{\coloneqq\;\tilde{x}_{0}^{\text{clean}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ (58)
where ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{T}W^{-1}\epsilon_{y}}$ is the extra noise introduced into $\tilde{x}_{0}^{\text{pm}}$ and will be further introduced into $x_{t-\Delta t}$:
$x_{t-\Delta t}$ $=a_{t}\tilde{x}_{0}^{\text{pm}}+b_{t}x_{t}+\epsilon_{\text{new}},$ (59)
@@ B.2 Addressing unknown noise and non-differentiable measurements
where $\epsilon_{\text{new}}$ the noise added by diffusion process, which should be specifically designed to ensure $x_{t-\Delta t}$ is sampled from correct distribution as in Eq. 3, i.e., the total noise $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. However, as $\epsilon_{y}$ is unknown noise, we have no information about the introduced noise $\epsilon_{\text{intro}}$. To solve this problem, we made the following principles: (i) despite that fact that we cannot guarantee $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$, we still hope $\epsilon_{\text{intro}}+\epsilon_{\text{new}}$ is as close to $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$ as possible; (ii) small $\eta_{1}$ and $\eta_{2}$ are helpful to maximize the alignment to measurement $y$. Notably,
@@ Appendix C Limitations
There are some limitations that deserve further study.
  • Theoretical extension. While MAS can, in principle, be generalized to nonlinear inverse problems, explicitly formulating the likelihood term $p(y\mid\tilde{x}_{0})$ becomes challenging. Developing effective sampling techniques under this setting is a promising direction for future research.
  • Improved prior modeling. Boys et al. [2] leverage Tweedie’s formula to compute the covariance of the Gaussian projection of $p(x_{0}\mid x_{t})$, which could be incorporated to enhance the MAS prior in Eq. 7a. Moreover, improved modeling of the likelihood $p(y\mid\tilde{x}_{0})$ may further boost the performance of MAS.
@@ Appendix E Pytorch-like code implementation
Here we provide a basic PyTorch-Like implementation of the calculation of $\tilde{x}_{0}^{\text{pm}}$ in Eq. 12, shown in LABEL:lst:mas_code.
1@torch.no_grad()
2def mas(
3 H, x0_hat, y,
4 eta_1=-0.2, eta_2=0.0
5):
6 bs, _, H_img, W_img = x0_hat.shape
7 x0_hat = x0_hat.view(bs, -1)
8 y = y.view( bs, -1) # measurement dim m
9 ut_y = H.Ut(y) # (bs, m)
10 singulars = H.singulars() # (m,)
11 nz = singulars > 0 # boolean mask
12 scale1 = 1.0 / (singulars[nz] ** 2 * eta_1 + eta_2)
13 ut_y[:, nz] = ut_y[:, nz] * scale1 # broadcasting OK
14 u_y = H.U(ut_y) # (bs, m)
15 rhs = x0_hat + H.Ht(u_y) # (bs, d)
16
17 vt_rhs = H.Vt(rhs) # (bs, d)
18 scale2 = 1.0 / (1.0 + singulars[nz]**2 / (singulars[nz]**2 * eta_1 + eta_2))
19 vt_rhs[:, nz] = vt_rhs[:, nz] * scale2
20 x0_pm = H.V(vt_rhs) # (bs, d)
21 x0_pm = x0_pm.view(bs, 3, H_img, W_img)
22 return x0_pm
Listing 1: PyTorch-like implementation of the calculation of $\tilde{x}_{0}^{\text{pm}}$ in Eq. 12.
@@ F.1 Details of the degradation operators
Super-resolution. We use the averagepooling downsampler as the forward operator.
@@
Figure 6: Restoration of inpaint (freeform) using MAS.
Inpaint (freeform). The freeform mask covers approximately $19.3\%$ of the total image area. Examples of restoration are shown in Fig. 6.
@@
Colorization. We simulate grayscale degradation by applying a fixed linear transformation to each pixel using the matrix $[\frac{1}{3},\frac{1}{3},\frac{1}{3}]$, replacing each RGB pixel with its average intensity.
@@ F.2 Details of the baseline models
$\Pi$GDM [22]. $\eta=1.0$, with DDIM sampler, as advised in the original paper.
@@
DAPS [27]. $\tau=0.01$, 100 steps of MCMC, with EDM sampler, as advised in the original paper.
@@ G.1 Computational time
The computational time of MAS on solving inverse problems is shown in Table 5. Our model achieves similar efficiency to DDNM and DDRM, demonstrating that MAS introduces minimal overhead while maintaining competitive runtime performance.
Table 5: Sampling time (Sec) per image of MAS on deblurring and super-resolution with FFHQ 256, evaluated using a single NVIDIA A6000 48G GPU. We set NFE=20 for all of the methods.
Method MAS $\Pi$GDM DDNM DDRM RED-Diff
@@
SR (8×8$\times$8 ×) 0.131 0.282 0.131 0.131 0.125
G.2 Comparison between posterior sampling and posterior mean
We show the comparison between posterior sampling Eq. 11 and posterior mean Eq. 12 in Fig. 7. The experiments results demonstrate that posterior mean achieve the best results. posterior sampling will introduce additional noise to the final restoration.
Figure 7: The comparison between posterior sampling in Eq. 11 and posterior mean Eq. 12 on deblurring task. We set $\eta_{1}=-0.2$ and $\eta_{2}=0$.
Table 6: Quantitative evaluation of solving super-resolution and deblurring. We set $\eta_{1}=-0.2$ and $\eta_{2}=0$ for all of the tasks.
Strength DDNM MAS
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
4$\times$ 36.77 0.080 0.930 64.60 36.71 0.075 0.929 43.81
8$\times$ 32.28 0.180 0.863 90.86 32.46 0.170 0.866 74.60
16$\times$ 28.29 0.282 0.800 107.95 28.50 0.272 0.805 101.52
32$\times$ 24.79 0.374 0.750 131.64 25.01 0.363 0.755 130.78
64$\times$ 21.69 0.449 0.720 152.05 21.86 0.441 0.724 151.10
128$\times$ 19.10 0.497 0.705 145.61 19.19 0.496 0.707 147.10
G.3 Noise-free inverse problem
Quantitative comparisons for solving noise-free inverse problems are presented in Table 7. The results demonstrate that our model consistently outperforms DDNM, achieving significantly lower FID scores.
We include additional quantitative evaluation of solving super-resolution and deblurring under different strength of degradation, see Table 8 and Table 9 for details. MAS constantly outperforms DDNM under different strength of degradation.
Table 7: Quantitative evaluation of solving noise-free linear inverse problems. We set $\eta_{1}=-0.2$, $\eta_{2}=0$ for all of the tasks.
Task Method NFE FFHQ ImageNet
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
SR 8×8$\times$8 × DDNM 20 32.28 0.180 0.863 90.76 28.13 0.294 0.693 150.42
DDRM 20 32.33 0.180 0.865 87.54 29.22 0.296 0.735 142.62
$\Pi$GDM 20 32.60 0.177 0.871 87.35 29.22 0.302 0.738 149.87
RED-Diff 20 31.26 0.286 0.823 160.28 28.64 0.362 0.714 254.99
MAS (ours) 20 32.46 0.170 0.866 76.80 29.11 0.282 0.732 135.17
DAPS 100 26.05 0.269 0.690 101.03 23.17 0.385 0.545 161.43
$\Pi$GDM 100 32.03 0.149 0.856 68.79 28.37 0.242 0.705 103.24
RED-Diff 100 33.52 0.191 0.887 113.99 29.87 0.311 0.752 157.61
Deblurring DDNM 20 41.60 0.063 0.967 44.82 37.37 0.107 0.912 51.50
DDRM 20 41.61 0.060 0.968 43.22 37.96 0.096 0.921 47.39
$\Pi$GDM 20 42.34 0.047 0.974 38.71 37.45 0.103 0.918 58.65
RED-Diff 20 23.82 0.446 0.665 185.93 23.13 0.501 0.560 314.05
MAS (ours) 20 42.34 0.041 0.971 33.27 38.25 0.085 0.918 37.09
DAPS 100 25.28 0.300 0.677 110.67 22.78 0.438 0.519 200.86
$\Pi$GDM 100 44.92 0.024 0.983 23.40 40.63 0.049 0.947 20.05
RED-Diff 100 28.51 0.339 0.814 167.57 26.47 0.507 0.676 302.68
Table 8: Quantitative evaluation of solving super-resolution of different strength of degradation. We set $\eta_{1}=-0.4$ and $\eta_{2}=0$ for all of the tasks.
Strength DDNM MAS
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
4$\times$ 36.77 0.080 0.930 64.60 36.71 0.075 0.929 43.81
8$\times$ 32.28 0.180 0.863 90.86 32.46 0.170 0.866 74.60
16$\times$ 28.29 0.282 0.800 107.95 28.50 0.272 0.805 101.52
32$\times$ 24.79 0.374 0.750 131.64 25.01 0.363 0.755 130.78
64$\times$ 21.69 0.449 0.720 152.05 21.86 0.441 0.724 151.10
128$\times$ 19.10 0.497 0.705 145.61 19.19 0.496 0.707 147.10
Table 9: Quantitative evaluation of solving deblurring of different strength of degradation. We set $\eta_{1}=-0.4$ and $\eta_{2}=0$ for all of the tasks.
Kernel size DDNM MAS
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
25 41.60 0.063 0.967 44.82 41.71 0.042 0.970 27.93
50 33.77 0.186 0.888 90.29 34.40 0.158 0.896 70.26
100 28.61 0.296 0.812 104.06 29.13 0.272 0.819 93.21
150 25.74 0.363 0.768 131.31 26.12 0.819 0.774 119.11
@@ G.3 Noise-free inverse problem
G.4 Adding Gaussian noise
We present the results of solving 8×8$\times$8 × super-resolution under varying levels of Gaussian noise in Fig. 9. The visualizations demonstrate that MAS maintains strong restoration performance, even under high noise conditions.
Figure 9: MAS for solving super-resolution (8×8$\times$8 ×) with various strength of Gaussian noise.
G.5 Adding non-differentiable degragation
We present the results of solving 8×8$\times$8 × super-resolution with non-differentiable degradations, including JPEG compression and quantization, in Fig. 10.

### v2

@@ B.1 Addressing Gaussian noise
Consider noisy image restoration problems in the form of $y=Hx+\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\epsilon_{y}}$, where $\epsilon_{y}$ is the added noise. Then the measurement $y$ can be decomposed to the sum of clean measurement $y^{\text{clean}}\coloneqq Hx$ and measurement noise $\epsilon_{y}$. Calculating $x_{0}^{*}$ leads to:
$x_{0}^{*}$ $=Y^{-1}[m_{0\mid t}+H^{\mathsf{T}}W^{-1}y]$ (25)
$=m_{0\mid t}+(Y^{-1}-\mathbb{I})m_{0\mid t}+Y^{-1}H^{T}W^{-1}y$ (26)
where $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{T}W^{-1}\epsilon_{y}$} is the extra noise introduced into $x_{0}^{*}$ and will be further introduced into $x_{t-\Delta t}$.
To address Gaussian noise with known variance, we modify Eq.˜7 and Eq.˜2 as:
$x_{0}^{*}=m_{0\mid t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}[(Y^{-1}-\mathbb{I})m_{0\mid t}+H^{\mathsf{T}}W^{-1}y]$ (27)
@@
Then $x_{0}^{*}$ is:
$x_{0}^{*}$ $=m_{0\mid t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}[(Y^{-1}-\mathbb{I})m_{0\mid t}+H^{\mathsf{T}}W^{-1}y]$ (29)
$=\underbrace{m_{0\mid t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}(Y^{-1}-\mathbb{I})m_{0\mid t}+Y^{-1}H^{T}W^{-1}y^{\text{clean}}}_{\coloneqq\;\tilde{x}_{0}^{\text{clean}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ (31)
@@
$x_{t-\Delta t}$ $=a_{t}x_{0}^{*}(t,x,y)+b_{t}x_{t}+\epsilon_{\text{new}},\quad\epsilon_{\text{new}}\sim\mathcal{N}(0,\Phi_{t})$ (32)
@@
Suppose $\Sigma_{t}=V\mathrm{diag}\{\lambda_{t1},\cdots,\lambda_{td}\}V^{T}$ $\Phi_{t}=V\mathrm{diag}\{\gamma_{t1},\cdots\gamma_{td}\}V^{T}$. Then the introduced noise $\epsilon_{\text{intro}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{t}\sigma_{y}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ is still a Gaussian distribution: $\epsilon_{\text{intro}}\sim\mathcal{N}(0,VD_{t}V^{T})$, with $D_{t}=diag\{d_{t1},\cdots,d_{td}\}$:
@@
The choice of and $\Phi_{t}$ need to ensure the total noise injected to $x_{t-\Delta t}$ conforms the iteration in Eq.˜2.
@@
Now we can yield $\epsilon_{\text{new}}$ by sampling from $\mathcal{N}(0,V\Gamma_{t}V^{T})$ to ensure that $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. We need to make sure $\lambda_{ti}$ guarantees the noise level of the introduced noise does not exceed the pre-defined noise level $c_{t}$, we also hope $\lambda_{t_{i}}$ as close as $1$ as possible. Therefore,
@@
In practice, we found that setting $\sigma_{y}$ slightly larger than the true $\sigma_{y}$ is beneficial, possibly because the denoiser is more sensitive to excessive noise.
B.2 Efficient calculation via SVD decomposition
Let $H=U\Sigma V^{\mathsf{T}}$ with singular values $s_{1},\dots,s_{n}$. Then
$W^{-1}=U\,\mathrm{diag}\!\Bigl(\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\Bigr)U^{\mathsf{T}},\qquad Y^{-1}=V\,\mathrm{diag}\!\Bigl(\tfrac{1}{\,1+s_{i}^{2}/(\eta_{1}s_{i}^{2}+\eta_{2})\,}\Bigr)V^{\mathsf{T}},$ (38)
see App.˜A for the proof. Hence both $W^{-1}v$ and $Y^{-1}u$ reduce to inexpensive diagonal scalings in the SVD basis, avoiding the calculation of any explicit matrix inversion or square-root. The algorithm of MAS for inverse problem is provided in Algorithm˜1.
As $\eta_{1}<0$, $W$ could be non-invertible. However, $W=U\text{diag}(\eta_{1}s_{1}^{2},\cdots,\eta_{1}s_{r}^{2},\eta_{2},\cdots,\eta_{2})U^{T}$. Hence $W$ is invertible if $\eta_{1}s_{i}^{2}+\eta_{2}\neq 0$ for every $i$. Even when $\eta_{2}=0$ and $\eta_{1}<0$ make $W$ singular, the update $W^{\dagger}y$ uses the Moore-Penrose pseudo-inverse $W^{\dagger}$, which is always well-defined. The pseudo‑inverse acts like an ordinary inverse on the range of $H$ and leaves the null‑space untouched, so the sampler remains stable. Empirically, small negative values ($0.5<\eta_{1}<0$) often give the visual boost without instability, as demonstrated in the ablation studies in Sec.˜5
B.3 Addressing unknown noise and non-differentiable measurements
As the measurement noise $\epsilon_{y}$ is non-Gaussian or unknown, it’s difficult to ensure the total noise introduced in $x_{t-\Delta t}$ is $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. In this case, we calculate $x_{0}^{*}$ using Eq.˜7 and update $x_{t-\Delta t}$ using Eq.˜2. Then $x_{0}^{*}$ is:
$x_{0}^{*}$ $=m_{0\mid t}+[(Y^{-1}-\mathbb{I})m_{0\mid t}+H^{\mathsf{T}}W^{-1}y]$ (39)
$=\underbrace{m_{0\mid t}+(Y^{-1}-\mathbb{I})m_{0\mid t}+Y^{-1}H^{T}W^{-1}y^{\text{clean}}}_{\coloneqq\;\tilde{x}_{0}^{\text{clean}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ (41)
where $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{T}W^{-1}\epsilon_{y}$} is the extra noise introduced into $x_{0}^{*}$ and will be further introduced into $x_{t-\Delta t}$:
$x_{t-\Delta t}$ $=a_{t}x_{0}^{*}+b_{t}x_{t}+\epsilon_{\text{new}},$ (42)
@@ B.3 Addressing unknown noise and non-differentiable measurements
where $\epsilon_{\text{new}}$ the noise added by diffusion process, which should be specifically designed to ensure $x_{t-\Delta t}$ is sampled from correct distribution as in Eq.˜2, i.e., the total noise $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. However, as $\epsilon_{y}$ is unknown noise, we have no information about the introduced noise $\epsilon_{\text{intro}}$. To solve this problem, we made the following principles: (i) despite that fact that we cannot guarantee $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$, we still hope $\epsilon_{\text{intro}}+\epsilon_{\text{new}}$ is as close to $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$ as possible; (ii) small $\eta_{1}$ and $\eta_{2}$ are helpful to maximize the alignment to measurement $y$. Notably,
@@ Appendix C Limitations
While MAS can, in principle, be generalized to nonlinear inverse problems, explicitly formulating the likelihood term $p(y\mid x_{\epsilon})$ becomes challenging. Developing effective sampling techniques under this setting is a promising direction for future research.
@@ Appendix E Pytorch-like code implementation
Here we provide a basic PyTorch-Like implementation of the calculation of $x_{0}^{*}$ in Eq.˜7, shown in Fig.˜6.
@torch.no_grad()
def mas(
H, x0_hat, y,
eta_1=-0.2, eta_2=0.0
bs, _, H_img, W_img = x0_hat.shape
x0_hat = x0_hat.view(bs, -1)
y = y.view( bs, -1) # measurement dim m
ut_y = H.Ut(y) # (bs, m)
singulars = H.singulars() # (m,)
nz = singulars > 0 # boolean mask
scale1 = 1.0 / (singulars[nz] ** 2 * eta_1 + eta_2)
ut_y[:, nz] = ut_y[:, nz] * scale1 # broadcasting OK
u_y = H.U(ut_y) # (bs, m)
rhs = x0_hat + H.Ht(u_y) # (bs, d)
[…]
scale2 = 1.0 / (1.0 + singulars[nz]**2 / (singulars[nz]**2 * eta_1 + eta_2))
vt_rhs[:, nz] = vt_rhs[:, nz] * scale2
x0_pm = H.V(vt_rhs) # (bs, d)
x0_pm = x0_pm.view(bs, 3, H_img, W_img)
return x0_pm
Figure 6: PyTorch-like implementation of the calculation of $x_{0}^{*}$ in Eq.˜7.
@@ F.1 Details of the degradation operators
Super-resolution. We use the downsampler with bicubic kernel as the forward operator.
@@
Inpaint (Random). Random Inpainting uses a generated random mask where each pixel has a 70% chance of being masked, following the settings in [26].
@@
Colorization. We simulate grayscale degradation by applying a fixed linear transformation to each pixel using the matrix $\frac{1}{3},\frac{1}{3},\frac{1}{3}]$, replacing each RGB pixel with its average intensity.
@@ F.2 Details of the baseline models
[…]
@@
DAPS [38]. $\tau=0.01$, with EDM sampler, as advised in the original paper.
@@ G.1 Computational time
The computational time of MAS on solving inverse problems is shown in Table˜5. Our model achieves similar efficiency to DDNM and DDRM, demonstrating that MAS introduces minimal overhead while maintaining competitive runtime performance.
Table 5: Sampling time (Sec) per image of MAS on deblurring and super-resolution with FFHQ 256, evaluated using a single NVIDIA A6000 48G GPU. We set NFE=20 and batch size = 20 for all of the methods.
[…]
@@
SR ($8\times$) 0.131 0.282 0.131 0.131 0.125
@@
G.2 Adding Gaussian noise
We present the results of solving $8\times$ super-resolution under varying levels of Gaussian noise in Fig.˜8. The visualizations demonstrate that MAS maintains strong restoration performance, even under high noise conditions.
Figure 8: MAS for solving super-resolution ($8\times$) with various strength of Gaussian noise.
G.3 Adding non-differentiable degragation
We present the results of solving $8\times$ super-resolution with non-differentiable degradations, including JPEG compression and quantization, in Fig.˜9.